OpenStudy (anonymous):

If instead of defining velocity as \(v(t) = x'(t)\) we had some other property of motion \(v(x) = t'(x)\) how would that alter things like acceleration and force?

OpenStudy (anonymous):

This property of motion would be in units of \(seconds/meter\)

OpenStudy (anonymous):

Maybe this would change the paradigm entirely. Just how arbitrary is it that we think about speed as meters/seconds, versus some other property of matter that would use units of seconds/meter.

OpenStudy (anonymous):

\[ t = t(x(t))\implies 1 = \frac{dt}{dx}\frac{dx}{dt}\implies t'(x) = \frac{1}{x'(t)} \]

OpenStudy (anonymous):

So\[ \frac{dx}{dt}= \frac{1}{\frac{dt}{dx}} \]Then \[ x''(t)=\frac{d^2x}{(dt)^2} = \frac{d}{dt}\left( \frac{1}{\frac{dt}{dx}}\right) \]

OpenStudy (anonymous):

\[ \frac{d}{dt}\left(\frac{1}{\frac{dt}{dx}}\right) = -\frac{\frac{d}{dt}\frac{dt}{dx}}{\frac{dt}{dx}\frac{dt}{dx}} \]

OpenStudy (anonymous):

What the heck would \[ \frac{d}{dt}\frac{dt}{dx} \]be?

OpenStudy (anonymous):

\[ \frac{dt}{dx} = t'(x) = t'(x(t)) \]

OpenStudy (anonymous):

\[ \frac{d}{dt}\frac{dt}{dx} = \frac{d^2t}{(dx)^2}\frac{dx}{dt} \]

OpenStudy (nincompoop):

why seconds/meter?

OpenStudy (anonymous):

Why meter/second?

OpenStudy (nincompoop):

at which something is accelerating or moving

OpenStudy (nincompoop):

you said, Just how arbitrary is it that we think about speed as meters/seconds, versus some other property of matter that would use units of seconds/meter.

OpenStudy (anonymous):

Well, seconds per meter is a more concrete than say \(f(x,t)\), which would be too abstract to example well.

OpenStudy (anonymous):

examine

OpenStudy (rational):

t(x) = x^3 + x what nonsense... time changes with respect to x ?

OpenStudy (nincompoop):

laughing out loud doggy

OpenStudy (rational):

lol how to think it ?

OpenStudy (anonymous):

I didn't say it was intuitive.

OpenStudy (anonymous):

Are you saying that time doesn't change?

OpenStudy (anonymous):

Because if time didn't change, then that would nothing could move.

OpenStudy (anonymous):

At the very least we can establish a relationship between time and distance.

OpenStudy (nincompoop):

so you're redefining position function as well?

OpenStudy (anonymous):

\(x\) is position. \(t\) is time.

OpenStudy (anonymous):

\(x(t)\) is position as a function of time. \(t(x)=x^{-1}(t)\), since \(t(x)\) is just time as a function of position.

OpenStudy (nincompoop):

how long was your physics class today? 20 meters

OpenStudy (anonymous):

Well, instead of doing a function inverse, you could also consider...\[ \overline{x}(t)=\frac{1}{x(t)} \]Maybe that would be another way to look at it.

OpenStudy (anonymous):

This is a bit less complicated. \[ \overline{v}(t)=\frac{d}{dt}\overline{x}(t)=\frac{d}{dt}\frac{1}{x(t)}=-\frac{x'(t)}{[x(t)]^2} \]

OpenStudy (nincompoop):

distance represents physical dimension, while time does not? 20-minute walk away 50-week ant sprint away 5-minute car drive is this what you're talking about ?

OpenStudy (anonymous):

The dimensional analysis would mean that \(\overline{v}(t)\) has units \[ \frac{m/s}{m^2} = \frac{1}{ms} \]

OpenStudy (anonymous):

The question here is... Suppose instead of thinking about the property velocity \(m/s\) we initially looked at motion differently, specifically as \(s/m\) or \(\Delta time / \Delta distance\). How would this affect other areas of physics?

OpenStudy (nincompoop):

so ... HOW FAR IS A second?

OpenStudy (anonymous):

Would there be a corresponding property to this property of motion similar how acceleration is to speed

OpenStudy (anonymous):

We are talking about a moving (but possibly at rest) object. There is a relationship between how it's time changes and it's motion changes.

OpenStudy (anonymous):

how its time changes and how its position changes

OpenStudy (nincompoop):

time can change independent of distance change distance can remain zero and time still continues to tick tock

OpenStudy (nincompoop):

dude stop trolling us laughing out loud

OpenStudy (anonymous):

I don't see how that matters here.

OpenStudy (anonymous):

Motion requires that time or distance change regardless.

OpenStudy (anonymous):

Quantum physics has a lot of intuitive stuff, even things being in multiple places at the same time.

OpenStudy (anonymous):

unintuitive stuff^

OpenStudy (anonymous):

So from a quantum physics point of view, saying that at any given time something has to have one position is somewhat absurd as well.